OpenStudy (anonymous):

write a quadratic equation with the given roots: 9-√11 and 9+√11

OpenStudy (***[isuru]***):

if u got a quadratic equation with root p , q u can write it as ( x - p)( x -q) and expand it.. in the same way u can write \[[x - ( 9 -\sqrt{11} ) ] [ x -(9 + \sqrt{11} ) ]\] as the quadratic equation u want and expand ........... can u expand this ?

OpenStudy (anonymous):

i keep getting the wrong answer x^2-18x-92 and x^2-18x+70

OpenStudy (anonymous):

come back!!!!

OpenStudy (***[isuru]***):

u cant get 2 expression.. u have to write a single quadratic expression with the given roots.. and u can get it by expanding the above expression.. can u expand it ?

OpenStudy (anonymous):

i know ive been getting single expressions but i cant get the right one can you give me the right answer? it might be my math that im doing wrong

OpenStudy (anonymous):

wait ive got it!!!! i was doing the math wrong. poitive 81 lol

OpenStudy (anonymous):

but since you are here can you help me with one that has the i's?

OpenStudy (***[isuru]***):

srry... i dont know much about complex numbers...

OpenStudy (anonymous):

1+√11i and 1-√11i

OpenStudy (anonymous):

oh ok thank you though

OpenStudy (anonymous):

\(i=\sqrt{-1}\)